# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

### v1

@@ IV-A EGM Driven Optimization
where $L\doteq log(I)$ is the logarithm of intensity, $p_{k}\in\{-1,+1\}$ indicates the polarity of brightness changes, and $t_{k}$ is the triggered timestamp. By leveraging the high temporal resolution of event cameras, we can effectively constrain the 3DGS optimization process in the temporal domain.
To incorporate the event data into the optimization process, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$. This allows us to form event frames at a higher temporal resolution than the original video frames:
@@
We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ between two video frames $I_{i}$ and $I_{i+1}$ using the accumulated events and the most recent frame:
@@
Then, the event-based rendering loss is introduced to enforce the consistency between rendered and latent intensity images:
@@
This loss function ensures that the 3DGS optimization process is constrained not only by the sparse video frames but also by the dense event data, leading to improved reconstruction quality. It is worth noting that when $j=0$, the latent intensity image $I_{t}$ is equal to the first most recent video frame $I_{i,0}$ and $\mathcal{L}_{EGM}$ reduces to the standard image rendering loss.
IV-B Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by the events, it does not explicitly exploit the motion information encoded in the event stream. To further improve the accuracy of camera poses and the reconstructed 3DGS, we introduce the Contrast Maximization (CMax) [28, 51, 52] framework and the Linear Event Generation Model (LEGM) [29, 30, 53] in this section. These models complement the EGM-driven optimization by extracting motion cues and gradient information from the event stream, providing additional constraints for the 3DGS optimization process.
Under the assumption of constant scene illumination, events are triggered solely by the motion of scene edges. The edges in the scene form continuous trajectories in the $(x,y,t)$ space. Intuitively, by back-projecting the events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the currently estimated motion. This insight motivates us to derive a motion field by leveraging the rendered depth from 3DGS using eq. 6 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the relative poses between adjacent frames.
Instead of warping each event as in the original CMax framework, we adopt a piece-wise warping approach for efficiency. Specifically, we first accumulate the events within each sub-interval to obtain event frames $E_{i,j}$ as in eq. 8 . Then, for randomly sampled timestamps $t_{ref}=t_{i,j}$, we warp the event frames in the neighboring $2r$ sub-intervals to $t_{ref}$:
$E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),$ (11)
where $m\in[-r,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow which can be derived from the rendered depth $\hat{D}$ in  eq. 6 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),$ (12)
@@ IV-B Unified CMax and LEGM Optimization
$\textrm{IPWE}_{i,j}=\frac{1}{2r+1}\sum_{m=j-r}^{j+r}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (13)
@@
Maximizing the contrast encourages the motion field to align with the actual motion of edges, leading to sharper IPWEs. This objective helps to improve the accuracy of the estimated camera poses and the geometric consistency of the 3DGS reconstruction. Furthermore, based on the Linear Event Generation Model (LEGM) [29, 30], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of a image):
@@
Based on this relationship, we formulate an additional gradient-based loss to constrain the 3DGS optimization:
$\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},$ (17)
@@
where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight coefficients for $\mathcal{L}_{cm}$ and $\mathcal{L}_{grad}$ respectively.
IV-C Photometric Bundle Adjustment
The aforementioned event-based constrains, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS. However, as event cameras only record brightness changes and lack color perception, directly applying them to 3DGS optimization may lead to inconsistent color rendering. To ensure cross-view consistency of the 3DGS rendering, we introduce the Photometric Bundle Adjustment (PBA) term.
Specifically, for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
@@ IV-C Photometric Bundle Adjustment
where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames. We select $\mathcal{F}$ to be the adjacent video frames in consideration of computation costs.
By minimizing $\mathcal{L}_{PBA}$ over sampled pixels in the event frames and neighboring video frames, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames.
Figure 3: The illustration of unified CMax and LEGM optimization. We warp nearby event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE). The byproduct IWPE is utilized to establish additional constrain on 3DGS in gradient domain.
IV-D Fixed-GS Training Strategy
As demonstrated in Fig. 10, the $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues. The absence of color information in event data poses a significant challenge for scene reconstruction in previous NeRF-based methods. NeRF models employ implicit representations [54], utilizing MLPs to store scene color and density information in a coupled manner. Consequently, color distortions can affect scene density estimations and vice versa. In contrast, 3DGS explicitly stores various attributes independently within Gaussian points. This characteristic allows us to address color distortion issues specifically while maintaining other attributes.
To leverage this advantage, we propose a two-stage Fixed-GS scene optimization strategy. In the first stage, optimization is performed across all timestamps:
$G_{\theta}^{*}=\underset{\alpha,r,s,f}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (20)
where $\alpha,r,s,f$ is the opacity, rotation, scale factor and spherical harmonics of Gaussians, and t is the sampled timestamp during training. This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions. The second stage focuses on recovering accurate color information. During this phase, optimization is conducted exclusively on video frames. Capitalizing on the explicit representation of Gaussians, we optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed.
[…]
IV-E Overall Training Pipeline
Assembling all loss terms, we get the overall loss function for events:
where $\lambda_{PBA}$ are the weighting factor. It is noteworthy that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We borrow the idea of dynamic scene allocation strategies from LocalRF [9]. Our overall training pipeline builds upon the pipeline of CF-3DGS [5] while introducing novel components to integrate event camera data for free-trajectory scene reconstruction. Further implementation details are provided in the Supplementary Material.
V Experiments
TABLE I: Quantitative Evaluations on Tanks and Temples. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
@@ V Experiments
TABLE II: Pose accuracy on Tanks and Temples. Note that we use COLMAP poses in Tanks and Temples as the “ground truth”. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods which require precomputed poses are excluded.
@@
Figure 4: Qualitative comparison for novel view synthesis on Tanks and Temples. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 5: Pose Estimation Comparison on Tanks and Temples. We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
V-A Dataset
V-A1 Tanks and Temples
We conduct comprehensive experiments on the Tanks and Temples dataset [55]. Similar to LocalRF [9], we adopt the same 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF [9], we spatially downsample the videos to a resolution of 1080p. Increasing motion speed makes frames further apart, hence it is effectively equivalent to reducing the frame rate. We employ various temporal downsamplings to evaluate the robustness of different methods to varying camera speed scenes. In the 6 FPS setting, we retain one in every five frames as in LocalRF [9]. We perform corresponding frame downsampling for the other frame rate settings (4 FPS, 3 FPS, 2 FPS, and 1 FPS). To synthesize realistic event data, we first upsample the original videos by [56] and then apply the simulator V2E [57].
V-A2 RealEv-DAVIS
Due to the current lack of datasets featuring free-trajectory captures from event cameras, we introduce the RealEv-DAVIS dataset, encompassing four outdoor scenes. We use the DAVIS346 camera, which simultaneously records image frames and events at a resolution of 346x260 pixels. To simulate real-world free-trajectory scenarios, we capture handheld footage at 25 FPS, with each sequence lasting 40 seconds. For SLOW scenarios, we downsample the video clips by keeping every second frame. We further reduce the frame rate in FAST scenarios by retaining only one frame out of every five.
V-B Implementation details
Our implementation is primarily based on the PyTorch framework, and we follow the optimization parameters by the configuration outlined in the 3DGS [4] unless otherwise specified. All the parameters are optimized using Adam [58] with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decay to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5, respectively, based on subsequent experiment results. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for following CMax warping. The intervals of neighboring warping $r$ in Cmax are set to 3. That is, we warp the three adjacent event frames to the current sampled timestamp.
V-C Experimental Setup
V-C1 Metrics
We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM [59], and LPIPS [60] to quantitatively assess the quality of the rendered images. To measure the accuracy of pose estimation, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [61, 62], as delineated in [10]. Given that $\text{RPE}_{r}$ and $\text{RPE}_{t}$ measure relative pose errors between adjacent frames, they are inherently influenced by the frame rate. To facilitate a fair comparison across different frame rate settings, we upsample the estimated poses to the same frame rate before evaluation. This enables a consistent comparison of relative pose estimation errors across varying frame rate configurations.
TABLE III: Quantitative Evaluations on RealEv-DAVIS. We select the top-performing methods from the previous evaluation. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
Methods Pose-Free Input SLOW FAST
hall building corner outdoor hall building corner outdoor

### v3

@@ 4.1 EGM Driven Optimization
where $L\doteq log(I)$ is the logarithm of intensity, $p_{k}\in\{-1,+1\}$ indicates the polarity of brightness changes, and $t_{k}$ is the triggered timestamp.
As shown in Fig. 2 (1.1), to leverage the high temporal resolution of events, we first divide the time interval between two adjacent video frames $I_{i}$ and $I_{i+1}$ into N smaller subintervals $\varepsilon_{i,j}=\{\mathbf{e}_{k}|t_{i,j}\leq t_{k}\leq t_{i,j+1},\Delta t=\frac{t_{i+1}-t_{i}}{N},t_{i,j}=t_{i}+j\cdot\Delta t\}$. This allows us to form accumulated event frames at a higher temporal resolution:
@@
We then reconstruct the latent intensity image $I_{t}$ at any intermediate time $t\in\{t_{i,j}\}$ by integrating the accumulated events with the most recent frame:
@@
This latent intensity image provides a supervisory signal for our event-based rendering loss:
@@
By enforcing consistency between rendered and latent intensity images, this loss effectively utilizes the brightness information encoded in event streams between adjacent frames, addressing the challenge of sparse viewpoints in high-speed scenarios.
4.2 Unified CMax and LEGM Optimization
While $\mathcal{L}_{EGM}$ leverages the brightness change information recorded by events, it does not explicitly exploit the motion information encoded in the event stream. To address this, we introduce the Contrast Maximization (CMax) [11, 34, 38, 15] framework and the Linear Event Generation Model (LEGM) [10, 13, 56]. These models complement the previous EGM-driven optimization.
Under constant scene illumination, events are triggered by the motion of scene edges, forming continuous trajectories in (x, y, t) space. As shown in Fig. 2 (1.2), by warping(back-projecting) events along the correct motion trajectories, we can obtain a sharp image of warped events (IWE). Therefore, the sharpness of the IWE can serve as an indication of the accuracy of the estimated motion. This insight motivates us to derive the motion field by leveraging the rendered depth from 3DGS using Eq. 2 and the relative camera motion between neighboring timestamps. By optimizing the sharpness of the IWE, we can obtain the optimal motion field, which in turn helps to improve the geometric accuracy of the 3DGS and the camera poses.
As shown in Fig. 3, for efficiency, we adopt a piece-wise warping approach instead of warping individual events. Specifically, for current timestamp $t_{ref}=t_{i,j}$, we warp the event frames from previous $r$ sub-intervals:
$\vspace{1pt}E_{i,j-m\rightarrow j}=\textrm{warp}(E_{i,j-m},F_{i,j\rightarrow j-m}),\vspace{1pt}$ (7)
where $m\in[0,r]$, $F_{i,j-m\rightarrow j}$ is the optical flow derived from the rendered depth $\hat{D}$ in  Eq. 2 and relative pose $T_{i,j\rightarrow j-m}$ between two timestamps:
$\vspace{1pt}F_{i,j\rightarrow j-m}=\Pi(T_{i,j\rightarrow j-m}\Pi^{-1}(x,y,\hat{D}))-(x,y),\vspace{1pt}$ (8)
@@ 4.2 Unified CMax and LEGM Optimization
$\textrm{IPWE}_{i,j}=\frac{1}{r+1}\sum_{m=j-r}^{j}E_{i,m\rightarrow j}\approx\frac{1}{C}\Delta L.$ (9)
@@
Furthermore, based on the Linear Event Generation Model (LEGM) [10, 13], the brightness change $\Delta L$ at pixel $\mathbf{u}$ can be approximated by the dot product of the image gradient $\nabla L$ and the optical flow $\dot{\mathbf{u}}$ (note that $L$ is the logarithm of an image):
@@
Note that to compute $F_{i,j\rightarrow j+1}$, we estimate $T_{i,j,j+1}$ by leveraging the assumption of locally linear motion from $T_{i,j,j-1}$ and $T_{i,j,j}$. Based on this relationship, we formulate an additional gradient-based loss:
$\vspace{1pt}\mathcal{L}_{grad}=||C\cdot\textrm{IPWE}_{i,j}-(\hat{L}(\mathbf{u})-\hat{L}(u+F_{i,j\rightarrow j+1}))||^{2},\vspace{1pt}$ (13)
@@
where $\lambda_{cm}$ and $\lambda_{grad}$ are the balancing weight.
4.3 Photometric Bundle Adjustment
The aforementioned event-based constraints, $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$, leverage the brightness change and motion information encoded in the event data to constrain 3DGS. However, as event cameras only record brightness changes and lack color perception, directly applying them to 3DGS optimization may lead to inconsistent color rendering. To ensure cross-view consistency of the 3DGS rendering, we introduce the Photometric Bundle Adjustment (PBA) term.
Specifically, as shown in Fig. 2 (1.3), for a randomly sampled timestamp $t\in\{t_{i,j}\}$, we establish the following photometric reprojection error:
@@ 4.3 Photometric Bundle Adjustment
where $\mathbf{u}^{\prime}=\Pi(T_{i,j-r\rightarrow j}\Pi^{-1}(x,y,\hat{D}(u)))$ represent the coordinate on target view projected from the pixel $\mathbf{u}$ of source view $I_{s}$, $\mathcal{P}$ denotes the pixel samples of current frame, and $\mathcal{F}$ is the candidates of target video frames. We select $\mathcal{F}$ to be the nearest previous video frame in consideration of computation costs.
By minimizing $\mathcal{L}_{PBA}$ across sampled views, we encourage the 3DGS model to produce geometrically and photometrically consistent renderings across events and video frames, thus effectively resolving color inconsistencies inherent in event data.
Figure 3: The illustration of unified CMax and LEGM optimization. We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE). The byproduct IPWE is utilized to establish additional constraints on 3DGS.
@@
Table 1: Quantitative evaluations on Tanks and Temples. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
@@
Table 2: Pose accuracy on Tanks and Temples. We use COLMAP poses in Tanks and Temples as the “ground truth”. The unit of $\textrm{RPE}_{r}$ is in degrees, ATE is in the ground truth scale and $\textrm{RPE}_{t}$ is scaled by 100. Those methods that require precomputed poses are excluded.
4.4 Fixed-GS Training Strategy
The $\mathcal{L}_{PBA}$ term alone is insufficient to fully mitigate color distortion issues. To further address this challenge, we propose a two-stage Fixed-GS scene optimization strategy that takes advantage of 3DGS’s explicit attribute representation. In the first stage, all the parameters are optimizable and the optimization is performed across all timestamps:
$G_{\theta}^{*},T_{i,j}^{*}=\underset{\mu,\alpha,r,s,f,T_{i,j}}{\operatorname{argmin}}~{}\mathcal{L}_{event},~{}t\in\{t_{i,j}\},$ (16)
where $\mu,\alpha,r,s,f$ is the position, opacity, rotation, scale factor and spherical harmonics of the Gaussians, and t is the sampled timestamp during training. This stage results in a scene reconstruction with accurate structure and brightness, albeit with potential color distortions due to the dominant colorless event supervision overwhelming the sparse RGB frame color supervision. The second stage focuses on recovering accurate color information. During this phase, optimization is conducted exclusively on video frames. We optimize only the spherical harmonic coefficients of the Gaussians while keeping other parameters fixed:
@@ 4.4 Fixed-GS Training Strategy
4.5 Overall Training Pipeline
Assembling all loss terms, we get the overall loss function:
where $\lambda_{PBA}$ are the weighting factor. Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We incorporate dynamic scene allocation strategies from LocalRF [28] for handling extended video sequences. Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS [9] while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction. Please refer to the Supplementary Material for the algorithm pipeline and additional implementation details.
5 Experiments
Figure 4: Qualitative comparison for novel view synthesis. The first two rows are from Tanks and Temples and the last row is from RealEv-DAVIS. Our approach produces more realistic rendering results with fine-grained details. Better viewed when zoomed in.
Figure 5: Pose estimation comparison. We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
5.1 Dataset
Tanks and Temples.  We conduct comprehensive experiments on the Tanks and Temples dataset [22]. Similar to LocalRF [28], we adopt 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF [28], we apply 4× spatial downsampling to the videos. To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS. The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios. To synthesize realistic event data, we first upsample the original videos by [18] and then apply the simulator V2E [14].
RealEv-DAVIS.  Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS, comprising four outdoor scenes. Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS. We employ COLMAP for ground-truth poses. For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames. Further details are provided in the supplementary material.
5.2 Implementation details
We follow the optimization parameters by the configuration outlined in the 3DGS [20]. We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decays to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for the following CMax warping. The intervals of neighboring warping $r$ in CMax are set to 3. The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS. We provide detailed ablation studies on these hyperparameters and additional implementation details in the supplementary material.
5.3 Experimental Setup
Metrics.  We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM [49], and LPIPS [54]. For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [55, 39], as delineated in [4]. Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
Baselines.  For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF [28] and F2-NeRF [48]. We also include pose-free methods like Nope-NeRF [4] and CF-3DGS [9]. (2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF [21], EvDeblurNeRF [5] and Event-3DGS [16]. Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens [43]) to temporally upsample frames before feeding them into CF-3DGS.
5.4 Experimental Results
We select every ten frames as a test image for NVS evaluation following LocalRF [28]. Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in iNeRF [53], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
Results on Tanks and Temples.  Tables 1 and 2 demonstrate two key findings: (1) Our event-aided approach achieves up to 3dB higher PSNR and nearly 40% lower trajectory error at 1FPS compared to frame-based methods, indicating the critical value of event data in high-speed scenarios. (2) Our method maintains 1.55dB PSNR advantage over Event-3DGS at 1FPS, confirming that our integration framework effectively exploits the nature of event data beyond merely using it. Fig. 4 shows our method produces sharper edges and finer textures, while Fig. 5 illustrates we achieve more accurate trajectory estimation.
Methods Input SLOW FAST
NVS Pose NVS Pose